# Transforming hyperspectral images into chemical maps: A new deep learning based approach to hyperspectral image processing

### v1

@@ Abstract
Current approaches to chemical map generation from hyperspectral images are based on models such as partial least squares (PLS) regression, generating pixel-wise predictions that do not consider spatial context and suffer from a high degree of noise. This study proposes an end-to-end deep learning approach using a modified version of U-Net and a custom loss function to directly obtain chemical maps from hyperspectral images, skipping all intermediate steps required for traditional pixel-wise analysis. We compare the U-Net with the traditional PLS regression on a real dataset of pork belly samples with associated mean fat reference values. The U-Net obtains a test set root mean squared error of between 9% and 13% lower than that of PLS regression on the task of mean fat prediction. At the same time, U-Net generates fine detail chemical maps where 99.91% of the variance is spatially correlated. Conversely, only 2.53% of the variance in the PLS-generated chemical maps is spatially correlated, indicating that each pixel-wise prediction is largely independent of neighboring pixels. Additionally, while the PLS-generated chemical maps contain predictions far beyond the physically possible range of 0-100%, U-Net learns to stay inside this range. Thus, the findings of this study indicate that U-Net is superior to PLS for chemical map generation.
@@ 1 Introduction
Hyperspectral imaging (HSI) is an advanced analytical technique that enables the acquisition of spatially distributed spectral information from material surfaces (Mäkelä et al. 33). Unlike traditional spectroscopy, which provides only bulk spectral data, HSI simultaneously captures spectral and spatial information, offering a more comprehensive analysis of heterogeneous samples.
@@
HSI is widely applied across various scientific disciplines and implemented in multiple spectral modalities, including near-infrared (NIR) (Albano-Gaglio et al. 3), Raman (Tao et al. 46), and infrared (IR) (Boseley et al. 14) spectroscopy. The fundamental distinction between HSI and conventional spectroscopy lies in its imaging capability, where each pixel of an image contains a complete spectral profile. This enables the generation of chemical maps representing the spatial distribution of specific chemical components within a sample. However, due to the high-dimensional nature of HSI data, advanced data modeling techniques are required to extract meaningful and interpretable information. The ability to visualize compositional variations in an image format makes HSI a powerful tool in diverse applications, including material science, pharmaceuticals, food quality assessment, and biomedical imaging.
@@
Chemometric modeling plays a crucial role in HSI data processing (Amigo et al. 5, Mäkelä et al. 33), particularly in the generation of chemical maps. Traditionally, chemical maps are generated using supervised or unsupervised methods, where each pixel is analyzed independently. For instance, supervised models are typically trained using mean spectra and subsequently applied to predict chemical properties for individual pixels (Amigo et al. 5). In the case of supervised modeling, it is essential to note that the minimum sample size required for chemical analysis is often larger than the pixel size in hyperspectral images. Consequently, acquiring pixel-wise reference values for chemical maps is generally infeasible. Instead, reference values are obtained as averaged measurements at a much lower resolution, such as over an entire sample or larger subsets thereof.
@@
Because reference values represent sample-wise means, it is common practice to train models to predict these mean values. A straightforward approach involves averaging the NIR hyperspectral image pixels to generate a mean spectrum, which is then paired with its corresponding mean reference value. These data pairs are typically used to train a Partial Least Squares (PLS) regression model (Wold 49, Wold et al. 50), which is subsequently applied to generate pixel-wise predictions, producing chemical maps (Thumm et al. 47, ElMasry et al. 17, 16, Albano-Gaglio et al. 2). However, this approach has several limitations. Traditional PLS-based chemical maps treat each pixel independently, resulting in predictions with a low degree of spatial correlation. Pixel-wise predictions may fall outside physically meaningful ranges (0-100 %), affecting interpretability. The lack of spatial structure limits the ability to assess variations within samples, thereby constraining downstream optimization.
@@
The significance of spatial and contextual information in HSI data modeling is now widely recognized in the chemometric domain (Jamme and Duponchel 29, Xu and Gowen 51). Recent studies have demonstrated the benefits of integrating spatial information alongside spectral data to enhance model accuracy (Gorretta et al. 24, Gaci et al. 23, Mishra et al. 36). These approaches offer improved pixel-level property prediction; however, they do not fundamentally enhance spatial coherence because the chemical maps are still generated in a pixel-wise manner.
@@
This study aims to bridge the gap between spectral-based models, such as PLS, and the inherently spatio-spectral task of chemical map generation. To this end, a novel method for chemical map generation is proposed. A modification of U-Net (Ronneberger et al. 39), a convolutional neural network (CNN) originally designed for semantic segmentation in medical imaging, is proposed for chemical map generation. In semantic segmentation, each pixel in an input image is to be classified into discrete categories. The equivalent task for chemical map generation involves continuous regression rather than classification. There are several advantages of using U-Net for chemical mapping. Unlike PLS, which processes pixels independently, U-Net jointly considers spatial and spectral information within hyperspectral images. It performs exceptionally well with limited labeled samples (Siddique et al. 41, Luo et al. 31, Azad et al. 9, Archana and Jeevaraj 8). Like PLS models trained on mean spectra, the proposed U-Net requires only mean reference values per sample. A custom multi-objective loss function trains U-Net to generate pixel-wise predictions strictly inside the physically valid range (0-100%) and leverages spatial structure to generate smoothly varying chemical maps, neither of which PLS accomplishes.
@@ 2.1.1 Dataset split
Albano-Gaglio et al. [2] split the dataset into two subsets, cross-validation (CV) and test, with the DUPLEX algorithm (Snee 43). The DUPLEX algorithm generates two datasets in an iterative process based on Mahalanobis distance (Mahalanobis 32) between the mean spectra of the pork bellies.
In the dataset split provided by Albano-Gaglio et al. [2], the CV set consists of 122 pork bellies, while the test set consists of 60 bellies. Albano-Gaglio et al. [2] randomly split the CV set into five folds (subsets). To make the results of this study comparable to those of Albano-Gaglio et al. [2], the same splitting of the CV set into folds from Albano-Gaglio et al. [2] was requested. However, as the randomly generated splits were not stored, this request could not be fulfilled. Instead, a modification of the DUPLEX algorithm to generate not two but six subsets of the dataset was explored. The original test set was kept as one of the subsets to enable a direct comparison between the test set results and those of Albano-Gaglio et al. [2]. The modified DUPLEX algorithm divides the CV set into five folds. Before feeding the mean spectra to DUPLEX, a dimensionality reduction using principal component analysis (PCA, Hotelling 27) was performed using six components, which was the minimum required to retain at least $99\%$ of the variance of the CV subset. In Table 1, a description of the number of pork bellies and how associated slices are distributed between the dataset splits yielded by DUPLEX is presented. In Fig. 3, distributions of reference fat percentages of these dataset splits are presented. All fat classes, F1-F5, are represented in the CV and test sets.
@@ 2.2.1 PLS
A PLS model using the pork bellies’ mean absorbance spectra was calibrated as a baseline. For a pork belly, the mean absorbance spectrum was calculated by first taking the negative logarithm of each pixel in each hyperspectral image of each pork belly slice and then computing the average across all pixels inside the slices’ masks. Subsequently, the spectra were preprocessed by application of Standard Normal Variate (SNV) transformation (Barnes et al. 12) followed by convolution with a Savitzky-Golay filter (Savitzky and Golay 40, Steinier et al. 45) with a window length of seven, a polynomial order of two, and a derivative order of two. These preprocessing steps combat Rayleigh scattering and reveal the chemical information in the spectra (Rinnan et al. 38, Sørensen et al. 44). Finally, a column-wise mean centering of the spectra and the reference fat values was performed using statistics computed only on the training sets.
A 5-fold cross-validation was used to determine the optimal number of components. For each CV fold, the root mean squared error (RMSE) on the validation set was computed for all possible numbers of components. Subsequently, the average over the five folds was calculated to get a mean RMSE for each number of components. This evaluation yielded nine components minimizing the RMSE, so a PLS on the entire CV set using nine components was recalibrated. Thus, the final PLS model was calibrated using all the available data points in the CV set. The PLS calibrations were carried out using the fast (Alin 4) and numerically stable (Andersson 6) Improved Kernel PLS Algorithm 1 (Dayal and MacGregor 15) with the software implemented by [20]. The CV was done with the same software using the fast algorithm by [22].
In addition to predicting from the mean spectra for each pork belly, this same PLS model was used to generate chemical maps by pixel-wise prediction as traditionally done in HSI processing (Amigo et al. 5, Mäkelä et al. 33). The same preprocessing was applied to each pixel as for the mean spectra described above.
@@ 2.2.2 U-Net
U-Net was modified in two ways: The first modification was replacing the upsampling by transposed convolution, as proposed by [39], with upsampling by bilinear interpolation. Transposed convolutions cause square-like artifacts in the output (Odena et al. 37). Bilinear interpolation avoids these artifacts while reducing the number of parameters in the model, thereby decreasing training time.
@@
A illustrates the model architecture and the sizes of intermediate tensors. Apart from these modifications, the new design followed the original U-Net implementation. As in the original U-Net implementation, no padding to any convolutions was applied, and only the so-called valid parts of the convolutions were evaluated. The valid parts of convolution are those pixels where the kernel completely overlaps with the input image or intermediate feature map. All convolution layers had their bias initialized to zero and their scale parameters initialized by sampling from a Kaiming He Normal Distribution (He et al. 25). While Ronneberger et al. [39] wrote the original U-Net implementation in Caffe (Vedaldi et al. 48), this study recreated it in PyTorch (Ansel et al. 7) to accommodate the customization needs better and to ensure easy compatibility and integration with other modern works.
@@
Similarly, the size of the spectral dimension was reduced by a factor of two from 124 channels to 62 channels. However, instead of strided convolutions, the spectral dimension was reduced by averaging neighboring wavelength channels. This reduction simultaneously reduces training time and alleviates the so-called curse of dimensionality, which is prevalent for deep learning on hyperspectral images (Signoroni et al. 42).
@@
Given a batch of hyperspectral images, let $\widehat{\mathbf{Y}}\in\mathbb{R}^{B\times H\times W}$ be the U-Net’s predicted chemical maps with batch size $B$, height $H$ pixels, and width $W$ pixels. For a given batch, consider also the associated reference fat values $\mathbf{y}\in\mathbb{R}^{B}$ and segmentation masks $\mathbf{M}\in\{0,1\}^{B\times H\times W}$. In $\mathbf{M}$, a value of 00 indicates background and a value of $1$ indicates foreground in the corresponding pixel of $\widehat{\mathbf{Y}}$. Such pixels are considered outside, respectively, inside the mask. For a batch of images, consider the modified U-Net’s masked prediction, $\widehat{\mathbf{Y}}^{\mathbf{M}}$, defined using element-wise multiplication ($\odot$)
@@
The true fat contents in each pixel must lie in the range $0-100\%$. However, only optimizing against the mean fat prediction provides no guarantee for a proper range of pixel-wise predictions. Thus, to aid the optimization process, the out-of-bounds loss (OOBL) was defined as the sum of squared deviations from the $0-100$ range in $\widehat{\mathbf{Y}}^{\mathbf{M}}$
@@
Here, the lambdas are scalar multipliers used to weigh the four loss terms relative to each other. This study used $\lambda_{MSE}=1$, $\lambda_{OOBL}=10^{-3}$, $\lambda_{SL}=20$, and $\lambda_{L2}=10^{-3}$.
To train the modified U-Net model using the loss function in Eq. (7), the Adam optimizer (Kingma and Ba 30) was used with an initial learning rate of $10^{-3}$, and hyperparameters $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The best model was chosen based on the validation MSE computed by Eq. (3) using the entire validation set as a single batch [^2] The whole validation set could not be stored on the GPU. Therefore, when encountering a sample, the squared errors were accumulated, returning the mean when all samples had been seen. This computation is mathematically equivalent to using the entire validation set as a batch. The model selection is based on MSE instead of TL, as this study’s primary interest is a solution minimizing MSE while using the other loss terms to guide the properties of such a solution (i.e., smoothly varying and inside the $0-100$ range). However, as evident in Section 3, MSE dominates the other terms in TL in practice. Therefore, using TL instead of MSE for model selection would likely yield similar results.
@@
The learning rate was left unchanged for the first 30 epochs, called the burn-in period. After the burn-in epochs, counters for early stopping and learning rate reduction were initiated. After 10 epochs without reaching a new set of best weights, the learning rate was reduced by a factor of 10 but not below $10^{-7}$. Immediately before lowering the learning rate, the model was restored to its current best weights, and the optimizer’s state was restored to its state at that time. If, after 30 epochs, no new set of best weights was achieved, training was halted, and the current best weights were restored and returned. If the early stopping was not triggered, the training was halted after 250 epochs. In practice, all training runs were halted by the early stopping criterion, not the 250 epoch limit.
As for PLS, the aim was to have a U-Net trained on all the available data in the CV set. This could be achieved for PLS by averaging the hyperparameter (number of components) over the CV folds and refit on the entire CV set using this hyperparameter setting. For U-Net, however, averaging hyperparameters such as early stopping epochs and epochs for learning rate reduction provide no guarantee that a U-Net trained with these hyperparameters will have its weights end up in a meaningful place. Instead, an ensemble was constructed from all five U-Nets to have a U-Net trained on the entire CV set. Then, the ensemble’s prediction is simply the uniform average of its constituents’ predictions. This ensemble U-Net will simultaneously have been trained and validated on all images in the CV set. Constructing such an ensemble has previously yielded good results for predicting protein content in NIR-HSI images of grain (Engstrøm et al. 19).

### v5

# Transforming Hyperspectral Images Into Chemical Maps: A Novel End-to-End Deep Learning Approach

@@ Abstract
Current approaches to chemical map generation from hyperspectral images are based on models such as partial least squares (PLS) regression, generating pixel-wise predictions that do not consider spatial context and suffer from a high degree of noise. This study proposes an end-to-end deep learning approach using a modified version of U-Net and a custom loss function to directly obtain chemical maps from hyperspectral images, skipping all intermediate steps required for traditional pixel-wise analysis. This study compares the U-Net with the traditional PLS regression on a real dataset of pork belly samples with associated mean fat reference values. The U-Net obtains a test set root mean squared error that is 7% lower than that of PLS regression on the task of mean fat prediction. At the same time, U-Net generates fine detail chemical maps where 99.91% of the variance is spatially correlated. Conversely, only 2.37% of the variance in the PLS-generated chemical maps is spatially correlated, indicating that each pixel-wise prediction is largely independent of neighboring pixels. Additionally, while the PLS-generated chemical maps contain predictions far beyond the physically possible range of 0-100%, U-Net learns to stay inside this range. Thus, the findings of this study indicate that U-Net is superior to PLS for chemical map generation.
@@ 1 Introduction
 [^1] This work is an update of the original work by Engstrøm et al. [18] and is the preferred version.  D describes the update in detail.
@@
Hyperspectral imaging (HSI) is an advanced analytical technique that enables the acquisition of spatially distributed spectral information from material surfaces (Mäkelä et al. 34). Unlike traditional spectroscopy, which provides only bulk spectral data, HSI simultaneously captures spectral and spatial information, offering a more comprehensive analysis of heterogeneous samples.
@@
HSI is widely applied across various scientific disciplines and implemented in multiple spectral modalities, including near-infrared (NIR) (Albano-Gaglio et al. 3), Raman (Tao et al. 48), and infrared (IR) (Boseley et al. 14) spectroscopy. The fundamental distinction between HSI and conventional spectroscopy lies in its imaging capability, where each pixel of an image contains a complete spectral profile. This enables the generation of chemical maps representing the spatial distribution of specific chemical components within a sample. However, due to the high-dimensional nature of HSI data, advanced data modeling techniques are required to extract meaningful and interpretable information. The ability to visualize compositional variations in an image format makes HSI a powerful tool in diverse applications, including material science, pharmaceuticals, food quality assessment, and biomedical imaging.
@@
Chemometric modeling plays a crucial role in HSI data processing (Amigo et al. 5, Mäkelä et al. 34), particularly in the generation of chemical maps. Traditionally, chemical maps are generated using supervised or unsupervised methods, where each pixel is analyzed independently. For instance, supervised models are typically trained using mean spectra and subsequently applied to predict chemical properties for individual pixels (Amigo et al. 5). In the case of supervised modeling, it is essential to note that the minimum sample size required for chemical analysis is often larger than the pixel size in hyperspectral images. Consequently, acquiring pixel-wise reference values for chemical maps is generally infeasible. Instead, reference values are obtained as averaged measurements at a much lower resolution, such as over an entire sample or larger subsets thereof.
@@
Because reference values represent sample-wise means, it is common practice to train models to predict these mean values. A straightforward approach involves averaging the NIR hyperspectral image pixels to generate a mean spectrum, which is then paired with its corresponding mean reference value. These data pairs are typically used to train a Partial Least Squares (PLS) regression model (Wold 51, Wold et al. 52), which is subsequently applied to generate pixel-wise predictions, producing chemical maps (Thumm et al. 49, ElMasry et al. 17, 16, Albano-Gaglio et al. 2). However, this approach has several limitations. Traditional PLS-based chemical maps treat each pixel independently, resulting in predictions with a low degree of spatial correlation. Pixel-wise predictions may fall outside physically meaningful ranges (0-100 %), affecting interpretability. The lack of spatial structure limits the ability to assess variations within samples, thereby constraining downstream optimization.
@@
The significance of spatial and contextual information in HSI data modeling is now widely recognized in the chemometric domain (Jamme and Duponchel 30, Xu and Gowen 53). Recent studies have demonstrated the benefits of integrating spatial information alongside spectral data to enhance model accuracy (Gorretta et al. 25, Gaci et al. 24, Mishra et al. 37). These approaches offer improved pixel-level property prediction; however, they do not fundamentally enhance spatial coherence because the chemical maps are still generated in a pixel-wise manner.
@@
This study aims to bridge the gap between spectral-based models, such as PLS, and the inherently spatio-spectral task of chemical map generation. To this end, a novel method for chemical map generation is proposed. A modification of U-Net (Ronneberger et al. 41), a convolutional neural network (CNN) originally designed for semantic segmentation in medical imaging, is proposed for chemical map generation. In semantic segmentation, each pixel in an input image is to be classified into discrete categories. The equivalent task for chemical map generation involves continuous regression rather than classification. There are several advantages of using U-Net for chemical mapping. Unlike PLS, which processes pixels independently, U-Net jointly considers spatial and spectral information within hyperspectral images. It performs exceptionally well with limited labeled samples (Siddique et al. 43, Luo et al. 32, Azad et al. 9, Archana and Jeevaraj 8). Like PLS models trained on mean spectra, the proposed U-Net requires only mean reference values per sample. A custom multi-objective loss function trains U-Net to generate pixel-wise predictions strictly inside the physically valid range (0-100%) and leverages spatial structure to generate smoothly varying chemical maps, neither of which PLS accomplishes.
@@ 2.1.1 Dataset split
Albano-Gaglio et al. [2] split the dataset into two subsets, cross-validation (CV) and test, with the DUPLEX algorithm (Snee 45). The DUPLEX algorithm generates two datasets in an iterative process based on Mahalanobis distance (Mahalanobis 33) between the mean spectra of the pork bellies.
In the dataset split provided by Albano-Gaglio et al. [2], the CV set consists of 122 pork bellies, while the test set consists of 60 bellies. Albano-Gaglio et al. [2] randomly split the CV set into five folds (subsets). To make the results of this study comparable to those of Albano-Gaglio et al. [2], the same splitting of the CV set into folds from Albano-Gaglio et al. [2] was requested. However, as the randomly generated splits were not stored, this request could not be fulfilled. Instead, a modification of the DUPLEX algorithm to generate not two but six subsets of the dataset was explored. The original test set was kept as one of the subsets to enable a direct comparison between the test set results and those of Albano-Gaglio et al. [2]. The modified DUPLEX algorithm divides the CV set into five folds. Before feeding the mean spectra to DUPLEX, a dimensionality reduction using principal component analysis (PCA, Pearson 39, Hotelling 28) was performed using six components, which was the minimum required to retain at least $99\%$ of the variance of the CV subset. In Table 1, a description of the number of pork bellies and how associated slices are distributed between the dataset splits yielded by DUPLEX is presented. In Fig. 3, distributions of reference fat percentages of these dataset splits are presented. All fat classes, F1-F5, are represented in the CV and test sets.
@@ 2.2.1 PLS
A PLS model using the pork bellies’ mean absorbance spectra was calibrated as a baseline. For a pork belly, the mean absorbance spectrum was calculated by first taking the negative logarithm of each pixel in each hyperspectral image of each pork belly slice and subsequently preprocessing the pixels by application of the Standard Normal Variate (SNV) transformation (Barnes et al. 12) followed by convolution with a Savitzky-Golay filter (Savitzky and Golay 42, Steinier et al. 47) with a window length of seven, a polynomial order of two, and a derivative order of two. Subsequently, a mean spectrum for each pork belly was computed by averaging all pixels inside the slices’ masks. The preprocessing steps combat Rayleigh scattering and reveal the chemical information in the spectra (Rinnan et al. 40, Sørensen et al. 46). Finally, a column-wise mean centering of the spectra and the reference fat values was performed using statistics computed only on the training sets.
A 5-fold cross-validation was used to determine the optimal number of components. For each CV fold, the root mean squared error (RMSE) on the validation set was computed for all possible numbers of components. Subsequently, the average over the five folds was calculated to get a mean RMSE for each number of components. This evaluation yielded 11 components minimizing the RMSE, so a PLS on the entire CV set using 11 components was recalibrated. Thus, the final PLS model was calibrated using all the available data points in the CV set. The PLS calibrations were carried out using the fast (Alin 4) and numerically stable (Andersson 6) Improved Kernel PLS Algorithm 1 (Dayal and MacGregor 15) with the software implemented by [21]. The CV was performed using the same software with the fast algorithm by [23].
In addition to predicting from the mean spectra for each pork belly, this same PLS model was used to generate chemical maps by pixel-wise prediction as traditionally done in HSI processing (Amigo et al. 5, Mäkelä et al. 34). The same preprocessing was applied to each pixel as for the mean spectra described above. Thus, PLS predictions on the mean spectra are equivalent to the average of pixel-wise predictions across chemical maps of slices of the same pork bellies.
@@ 2.2.2 U-Net
U-Net was modified in two ways: The first modification was replacing the upsampling by transposed convolution, as proposed by [41], with upsampling by bilinear interpolation. Transposed convolutions cause square-like artifacts in the output (Odena et al. 38). Bilinear interpolation avoids these artifacts while reducing the number of parameters in the model, thereby decreasing training time.
@@
A illustrates the model architecture and the sizes of intermediate tensors. Apart from these modifications, the new design followed the original U-Net implementation. As in the original U-Net implementation, no padding to any convolutions was applied, and only the so-called valid parts of the convolutions were evaluated. The valid parts of convolution are those pixels where the kernel completely overlaps with the input image or intermediate feature map. All convolution layers had their bias initialized to zero and their scale parameters initialized by sampling from a Kaiming He Normal Distribution (He et al. 26). While Ronneberger et al. [41] wrote the original U-Net implementation in Caffe (Vedaldi et al. 50), this study recreated it in PyTorch (Ansel et al. 7) to accommodate the customization needs better and to ensure easy compatibility and integration with other modern works.
@@
Similarly, the size of the spectral dimension was reduced by a factor of two from 124 channels to 62 channels. However, instead of strided convolutions, the spectral dimension was reduced by averaging neighboring wavelength channels. This reduction simultaneously reduces training time and alleviates the so-called curse of dimensionality, which is prevalent for deep learning on hyperspectral images (Signoroni et al. 44).
@@
Given a batch of hyperspectral images, let $\widehat{\mathbf{Y}}\in\mathbb{R}^{B\times H\times W}$ be the U-Net’s predicted chemical maps with batch size $B$, height $H$ pixels, and width $W$ pixels. For a given batch, consider also the associated reference fat values $\mathbf{y}\in\mathbb{R}^{B}$ and segmentation masks $\mathbf{M}\in\{0,1\}^{B\times H\times W}$. In $\mathbf{M}$, a value of 0 indicates background and a value of $1$ indicates foreground in the corresponding pixel of $\widehat{\mathbf{Y}}$. Such pixels are considered outside, respectively, inside the mask. For a batch of images, consider the modified U-Net’s masked prediction, $\widehat{\mathbf{Y}}^{\mathbf{M}}$, defined using element-wise multiplication ($\odot$)
@@
The true fat contents in each pixel must lie in the range $0-100\%$. However, only optimizing against the mean fat prediction provides no guarantee for a proper range of pixel-wise predictions. Thus, to aid the optimization process, the out-of-bounds loss (OOBL) was defined as the sum of squared deviations from the 01000-100 range in $\widehat{\mathbf{Y}}^{\mathbf{M}}$
@@
Here, the lambdas are scalar multipliers used to weigh the four loss terms relative to each other. This study used $\lambda_{MSE}=1$, $\lambda_{OOBL}=10^{-3}$, $\lambda_{SL}=20$, and $\lambda_{L2}=10^{-3}$. These values were chosen to ensure that MSE was the numerically dominant term while the other terms still made up a non-negligible part of the loss function. In practice, the other terms were between one and two orders of magnitude lower than the MSE. This ensures that predicting an accurate mean fat content remains the primary optimization objective.
To train the modified U-Net model using the loss function in Eq. (7), the Adam optimizer (Kingma and Ba 31) was used with an initial learning rate of $10^{-3}$, and hyperparameters $\beta_{1}=0.9$, and $\beta_{2}=0.999$. The best model was chosen based on the validation MSE computed by Eq. (3) using the entire validation set as a single batch.  [^3] The whole validation set could not be stored on the GPU. Therefore, when encountering a sample, the squared errors were accumulated, returning the mean when all samples had been seen. This computation is mathematically equivalent to using the entire validation set as a batch. The model selection is based on MSE instead of TL, as this study’s primary interest is a solution minimizing MSE while using the other loss terms to guide the properties of such a solution (i.e., smoothly varying and inside the 01000-100 range). However, as evident in Section 3, MSE dominates the other terms in TL in practice. Therefore, using TL instead of MSE for model selection would likely yield similar results.
@@
The learning rate was left unchanged for the first 30 epochs, called the burn-in period. After the burn-in period, counters for early stopping and learning rate reduction were initiated. After 10 epochs without reaching a new set of best weights, the learning rate was reduced by a factor of 10 but not below $10^{-7}$. Immediately before lowering the learning rate, the model was restored to its current best weights, and the optimizer’s state was restored to its state at that time. If, after 30 epochs, no new set of best weights was achieved, training was halted, and the current best weights were restored and returned. If the early stopping was not triggered, the training was halted after 250 epochs. In practice, all training runs were halted by the early stopping criterion, not the 250 epoch limit.
As for PLS, the aim was to have a U-Net trained on all the available data in the CV set. This could be achieved for PLS by averaging the hyperparameter (number of components) over the CV folds and refit on the entire CV set using this hyperparameter setting. For U-Net, however, averaging hyperparameters such as early stopping epochs and epochs for learning rate reduction provide no guarantee that a U-Net trained with these hyperparameters will have its weights end up in a meaningful place. Instead, an ensemble was constructed from all five U-Nets to have a U-Net trained on the entire CV set. Then, the ensemble’s prediction is simply the uniform average of its constituents’ predictions. This ensemble U-Net will simultaneously have been trained and validated on all images in the CV set. Constructing such an ensemble has previously yielded good results for predicting protein content in NIR-HSI images of grain (Engstrøm et al. 20).